# XtalOpt Version 14: Variable-Composition Crystal Structure Search for Functional Materials Through Pareto Optimization

### v2

@@ 1 Introduction
Various global optimization strategies have been implemented within commonly used software platforms for CSP, such as evolutionary algorithms (XtalOpt [26, 16], MAISE [27], GASP [28], and USPEX [29]), particle swarm optimization (CALYPSO [30]), and random structure searching (AIRSS [31]). Further extension of the basic search algorithms to multi-objective global optimization (MOGO) approaches [32, 33] have facilitated the ground state search for functional materials [34, 35, 36, 37, 38, 39, 40, 41, 42, 43].
@@
XtalOpt, an evolutionary algorithm designed for CSP, is a fully open-source and cross-platform software package, available in both command-line interface (CLI) and graphical user interface (GUI) modes. It is capable of performing both single-objective (SO) and multi-objective (MO) searches [59] for the ground state structures of crystalline systems. Originally designed to search over a fixed chemical composition, previous implementations of XtalOpt [26, 60, 61, 62, 16, 59] start by producing an initial set of structures generated from specified supercells of the given empirical formula. Following successful local optimization of the population members, the selection of parent structures was based on the fitness, determined from the normalized relative enthalpies obtained in an SO search, or normalized relative value of user-defined objectives in a MO search [43, 62, 63, 64]. The offspring would be created from the parent structure(s), with the size and composition matching those of the initial user-defined list, through application of genetic operations. These genetic operations included: “crossover” where two parent structure are mixed to generate an offspring with the same composition, as well as “stripple” and “permustrain” mutations that generate a new structure by applying random distortions to a selected parent structure. Filtration of the parent pool was introduced to keep only those structures with desirable characteristics [65].
@@
Figure 1: The workflow in XtalOpt version 14, which offers a variable composition multi objective (VC-MO) search. Post relaxation, each structure’s energies are used to obtain it’s distance above the convex hull, which is combined with the values of the user-specified objectives to determine the structure’s suitability for parenting new offspring. Genetic operations are then applied to the selected parents to create new structures with compositions determined by the search type.
@@
The details of the implemented methodologies and instructions for using them is discussed in the following sections. While most of the described options are available in version 14.0 of XtalOpt, the multi-cut crossover functionality (Section 5.1) and the option to set a minimum total atom counts (Section 4) are introduced in versions 14.1 and 14.2, respectively.
@@ 2 Input chemical formulas
In the new version of XtalOpt, introducing the chemical system and (optionally) desired formula units differs in the following ways from the previous versions: (i) the “empiricalFormula” and “formulaUnits” pair of flags are removed, and (ii) the chemical system is now entirely specified by the entries of a single input flag “chemicalFormulas”.
@@
    Must include the “full chemical formula” (i.e., combination of element symbols and corresponding atom counts, e.g., “Ti1O4” instead of “TiO4”).
That is to say the input:
@@
The “chemicalFormulas” parameter not only determines the initial cell(s) to be created in the first generation, it also has an important role in specifying what type of the evolutionary search should be conducted; as detailed in the next section.
@@ 3 Various types of evolutionary searches
Generally, the overall search workflow of the new version of XtalOpt is similar to the previous versions: an initial set of structures is generated from the compositions and cell sizes specified in the input, and the genetic operations are applied to the selected parents from the pool to produce offspring. However, now, the details of how the genetic operations make new offspring and the children that can be created are different. That is to say, depending on the search type, the size and composition of the offspring produced by genetic operations can be different than that of their parent structure(s) or the input chemical formula list. The search type is determined by the details of the input formulas list and a set of new flags, as follows.
@@ 3.2 Multi-composition search
This input instructs XtalOpt to perform a “multi-composition” (MC) search. This type of search is similar to the FC search in that the initial set of structures are generated as usual, and subsequent individuals are forced to have a composition matching “one of the compositions specified in the input list”. Since the initial list now covers various compositions, however, the parent structure(s) can be chosen from different compositions. This extends “desired” motifs found in one composition to other ones, potentially speeding up the search in finding the best candidate structures for various compositions.
@@ 5 Genetic operations
In XtalOpt version 14, the crossover operation is extended by allowing for “multi-cut” merging of the parent cells, and two new genetic operations (applicable only to the VC search) and a new random mutation (relevant to all types of evolutionary searches), are introduced. Moreover, the input entries defining the probability of choosing the various genetic operations and the interpretation of their values have been changed, as follows.
@@ 5.1 Multi-cut crossover
Prior to XtalOpt version 14, the crossover operation was performed by a single-point “cut and splice” of the parent structures [26]. The user now has the option to instruct the code to cut the parent unit cells in multiple points to produce “ribbons” by using the input flag:
@@
Generally, if this flag is set to “$n$”, the code will cut each parent structure in “$n$+1” ribbons, and will select the atoms of the offspring structure from the alternative ribbons of the parent structures. The width of these ribbons is initialized equally, and then distorted randomly. For the case of the single-cut crossover, however, the setting for the minimum contribution of each parent is applied. By default, XtalOpt will perform a single-cut crossover, unless the user instructs otherwise. This setting can be modified during the run. It should be noted that specifying an “even” number of cut points will result in an “odd” number of ribbons, hence, an uneven selection of atoms for the produced structure.
@@ 5.2 New genetic operation: permutomic
To ensure full sampling of the stoichiometry space, a new evolutionary operation, “permutomic”, has been introduced in the VC search. This operation randomly adds (removes) an atom to (from) a structure chosen from the parent pool as schematically illustrated in Fig. 2(a). For such a search, the user should provide a probability for performing this operation in the input. This probability is adjustable during the run.
@@ 5.4 Chances of performing the genetic operations
In XtalOpt version 14, instead of using a “percent chance” for applying the various genetic operations, the probability of choosing a particular operation should be specified by a “relative weight”.
In the CLI input the relative weights can be specified using the following set of flags (instead of the “percentChances…” flag):
@@
Assuming that the set of relative weights of the genetic operations are given by $\{P_{i}\}$, XtalOpt determines the percent chance of applying the $j^{\textrm{th}}$ genetic operation, $C_{j}$, as:
$C_{j}=100\ \times\ \frac{P_{j}}{\sum\limits_{i\textrm{ = all relevant operations}}P_{i}}$
@@
Table 1: Input flags to specify the relative weights of various genetic operations, with their default values. The actual percentage chances of applying a genetic operation is listed in the last columns for the VC and FC/MC search types.
@@ 5.5 Secondary mutation: random supercell expansion
If this flag is specified to be a non-zero value, the expansion is performed by randomly choosing three integers as multiples for the lattice vectors, such that the produced supercell has up to “maxAtoms” total atoms in the cell. Further, a randomly chosen atom in the supercell is displaced randomly, subject to the minimum interatomic distances settings (Fig. 3). This operation can be useful in ground state search of systems where charge density waves lower the energy of the system.
@@ 6 Reference energies
As of version 14, XtalOpt uses the distance above the convex hull as the target value for the global energy minimization, as first introduced by Zunger and co-workers [44]. The calculation of the convex hull (performed using the Qhull library [67]), requires defining the reference energies. By default, the code uses “0.0” as the “elemental reference energies”. To obtain the correct convex hull for MC and VC searches, however, the correct reference energies need to be specified.
@@
The energies introduced in the input must have the same unit that is used by the local optimizer (e.g. meV, eV or Hartree) for the chemical formula listed to produce the correct results.
@@ 7 Pareto optimization and parent selection
In release 13 of XtalOpt [59] a basic generalized fitness function, calculated as the weighted sum of normalized values of objectives, was introduced to perform a MOGO search for materials with desired functionalities. By default XtalOpt employs this scalar fitness function for global optimization (i.e., the “Basic” optimization type) regardless of the number of objectives. In the new version of XtalOpt Pareto optimization, which is in principle applicable to both SO and MO searches, has been implemented as an alternative (the “Pareto” optimization type). The user can instruct the code to utilize Pareto optimization through the input flag:
@@
The Pareto-based fitness value relies on the calculated rank of structure, $r$, which is a value in the $[0,N-1]$ range , where $N$ is the total number of Pareto fronts and $r=0$ corresponds to the global Pareto front. Using the obtained ranks, solutions of each front are assigned a raw fitness value of
$f^{0}_{r}=\frac{N-r}{N}$ (1)
that corresponds to a fitness of $1$ for structures on the global Pareto front and $\frac{1}{N}$ for those on the last Pareto front. Next, the calculated crowding distances of structures in each front are scaled to the $[0.1,1]$ range, where the non-zero lower limit is chosen to eliminate the overlap between solutions of two successive fronts.
Finally, the Pareto-based fitness for the $i^{th}$ structure in the $r^{th}$ Pareto front with a scaled crowding distance of $d_{r,i}$ is calculated using the following formula:
$f_{r,i}=f^{0}_{r}-\left(\frac{1-d_{r,i}}{N}\right)\equiv f^{0}_{r+1}+\frac{d_{r,i}}{N}$ (2)
The above procedure is designed to: (i) ensure the rank-precedence by introducing a gap of at least $\frac{0.1}{N}$ between the “worst” solution of rank $r$ and the “best” solution of rank $r+1$, (ii) diversify the selection pool by prioritizing more unique candidate structures in each rank through applying the crowding distance values, and (iii) obtain a scalar fitness value in the range of $[0,1]$.
@@
which results in performing tournament selection only using the structural ranking, or calculating Pareto-based scalar fitness with $d_{r,i}=1$ for all structures (i.e., the raw fitness values obtained from Eq. 1).
@@
where the default value of “$-1$” implies not rounding the objective values.
@@ 9 Similarity check
The first version of XtalOpt used an approximate scheme, identifying duplicates by comparing their spacegroups, volumes and enthalpies [26] to within a user specified tolerance. Because this method was prone to false positives, an exact comparison technique that searches for transformations mapping atoms from one structure to another was developed. The resulting XtalComp library [69] was used to find duplicates since versions 9 [61] of XtalOpt. Duplicates were marked as such in the output and removed from the run. Though this exact matching scheme was key for finding certain inorganic compounds [70], it is dependent upon the choice of the user-specified tolerance and does not work well for molecular systems. For example, it would likely fail to flag structures that differ by a slight rotation of a molecule or a functional group as duplicates because the required tolerance would be too large.
@@
By default, the RDF tolerance is set to zero; hence XtalComp is employed for duplicate checking.
@@ 10 Seed structures
One example where providing sub-system seeds could be useful is for obtaining energies of reference structures, to be subsequently used in the convex hull construction, that are consistent with the local optimization settings of the search. This option, however, should be utilized carefully since a failure of the local relaxation of the reference structure by the external code will result in an incorrect convex hull leading to errors in fitness determinations. Introduction of sub-systems may also be useful if they contain certain structural motifs that could be present in the full system.
The seed structures, in general, are not required to comply with the minimum and maximum number of atom settings. Once a seed structure is read in with a total atom counts that is not consistent with the existing settings, the atom counts limits will be updated, while a warning message is produced in the output.
@@ 11 Construction of molecular units
For a MC or VC search, it is possible to use molecular units for the construction of the initial random generation of structures, as first introduced in XtalOpt version 10 [74] for a FC search. However, now the number of atoms of each type that can participate in the construction of molecular units is limited to the smallest number of atoms of that type in the input list of chemical formulas, regardless of search type.
@@ 12 Interface to machine-learning interatomic potentials
An easy-to-use Python script, vasp_uip.py, is included in the new release of XtalOpt to facilitate such calculations. This wrapper can employ a series of the recently developed universal interatomic potentials (UIPs), MACE [52] and CHGNet [53], Orb [54], and MatterSim [55], to perform local structure optimization using the Atomic Simulation Environment (ASE) library [75]. The input/output format is that of the VASP code, i.e., the script reads in a POSCAR file and produces a CONTCAR and a minimal OUTCAR file, which follows the VASP format for the outputted entries.
@@
With the above-mentioned input and output file formats, this script is designed to be used in the XtalOpt runs as a VASP optimizer, allowing for the efficient ground state search for a desired chemical system. It should be noted that the implemented potentials are chosen as prototypes of the increasingly popular UIPs, and extending the script to support other similar platforms is straightforward.
@@ 13.1 Input entry conventions
In general, and for consistency, the input pattern for the CLI flags are designed with the following conventions:

### v3

@@ 1 Introduction
Various global optimization strategies have been implemented within commonly used software platforms for CSP, such as evolutionary algorithms (XtalOpt [26, 16], MAISE [27], GASP [28], and USPEX [29]), particle swarm optimization (CALYPSO [30]), and random structure searching (AIRSS [31]). We have previously illustrated that the commonly used evolutionary algorithm is likely to have a higher success rate of finding the ground state structure in fixed-composition searches as compared to a purely random search [16]. However, employing constraints based upon chemical and physical intuition has made random structure searching (e.g., as implemented in the AIRSS code) a very powerful technique. Further extension of the basic search algorithms to multi-objective global optimization (MOGO) approaches [32, 33] has facilitated the ground state search for functional materials [34, 35, 36, 37, 38, 39, 40, 41, 42, 43].
@@
XtalOpt, an evolutionary algorithm designed for CSP, is a fully open-source and cross-platform software package, available in both command-line interface (CLI) and graphical user interface (GUI) modes. It is capable of performing both single-objective (SO) and multi-objective (MO) searches [59] for the ground state structures of crystalline systems. Originally designed to search over a fixed chemical composition, previous implementations of XtalOpt [26, 60, 61, 62, 16, 59] start by producing an initial set of structures generated from specified supercells of the given empirical formula. Following successful local optimization of the population members, the selection of parent structures was based on the fitness, determined from the normalized relative enthalpies obtained in an SO search, or normalized relative value of user-defined objectives in a MO search [43, 62, 63, 64]. The offspring would be created from the parent structure(s), with the size and composition matching those of the initial user-defined list, through application of genetic operations. These genetic operations included: “crossover” where two parent structures are mixed to generate an offspring with the same composition, as well as “stripple” and “permustrain” mutations that generate a new structure by applying random distortions to a selected parent structure. Filtration of the parent pool was introduced to keep only those structures with desirable characteristics [65].
@@
Figure 1: The workflow in XtalOpt version 14, which offers a variable-composition multi-objective (VC-MO) search. Post relaxation, each structure’s energies are used to obtain its distance above the convex hull, which is combined with the values of the user-specified objectives to determine the structure’s suitability for parenting new offspring. Genetic operations are then applied to the selected parents to create new structures with compositions determined by the search type.
@@
The details of the implemented methodologies and instructions for using them are discussed in the following sections. While most of the described options are available in version 14.0 of XtalOpt, the multi-cut crossover functionality (Section 5.1) and the option to set a minimum total atom count (Section 4) are introduced in versions 14.1 and 14.2, respectively.
@@ 2 Input chemical formulas
In the new version of XtalOpt, introducing the chemical system and (optionally) desired formula units differs in the following ways from the previous versions: (i) the “empiricalFormula” and “formulaUnits” pair of flags is removed, and (ii) the chemical system is now entirely specified by the entries of a single input flag “chemicalFormulas”.
@@
    Must include the “full chemical formula” (i.e., a combination of element symbols and corresponding atom counts, e.g., “Ti1O4” instead of “TiO4”).
That is to say, the input:
@@
The “chemicalFormulas” parameter not only determines the initial cell(s) to be created in the first generation, it also has an important role in specifying what type of the evolutionary search should be conducted, as detailed in the next section.
@@ 3 Various types of evolutionary searches
Generally, the overall search workflow of the new version of XtalOpt is similar to the previous versions: an initial set of structures is generated from the compositions and cell sizes specified in the input, and the genetic operations are applied to the selected parents from the pool to produce offspring. However, now, the details of how the genetic operations make new offspring and the children that can be created are different. That is to say, depending on the search type, the size and composition of the offspring produced by genetic operations can be different than that of their parent structure(s) or the input chemical formula list. The search type is determined by the details of the input formulas list and a set of new flags, as described in the next subsections.
@@ 3.2 Multi-composition search
This input instructs XtalOpt to perform a “multi-composition” (MC) search. This type of search is similar to the FC search in that the initial set of structures is generated as usual, and subsequent individuals are forced to have a composition matching “one of the compositions specified in the input list”. Since the initial list now covers various compositions, however, the parent structure(s) can be chosen from different compositions. This extends “desired” motifs found in one composition to other ones, potentially speeding up the search in finding the best candidate structures for various compositions.
@@ 5 Genetic operations
In XtalOpt version 14, the crossover operation is extended by allowing for “multi-cut” merging of the parent cells, and two new genetic operations (applicable only to the VC search) and a new random mutation (relevant to all types of evolutionary searches), are introduced. Moreover, the input entries defining the probability of choosing the various genetic operations and the interpretation of their values have been changed, as described below.
@@ 5.1 Multi-cut crossover
Prior to XtalOpt version 14, the crossover operation was performed by a single-point “cut and splice” of the parent structures [26]. The user now has the option to instruct the code to cut the parent unit cells at multiple points to produce “ribbons” by using the input flag:
@@
Generally, if this flag is set to “nn”, the code will cut each parent structure in “nn+1” ribbons, and will select the atoms of the offspring structure from the alternating ribbons of the parent structures. The width of these ribbons is initialized equally, and then distorted randomly. For the case of the single-cut crossover, however, the setting for the minimum contribution of each parent is applied. By default, XtalOpt will perform a single-cut crossover, unless the user instructs otherwise. This setting can be modified during the run. It should be noted that specifying an “even” number of cut points will result in an “odd” number of ribbons, hence, an uneven selection of atoms for the produced structure.
@@ 5.2 New genetic operation: permutomic
To ensure full sampling of the stoichiometry space, a new evolutionary operation, “permutomic”, has been introduced in the VC search. This operation randomly adds (removes) an atom to (from) a structure chosen from the parent pool, as schematically illustrated in Fig. 2(a). For such a search, the user should provide a probability for performing this operation in the input. This probability is adjustable during the run.
@@ 5.4 Chances of performing the genetic operations
In XtalOpt version 14, instead of using a “percent chance” for applying the various genetic operations, the probability of choosing a particular operation should be specified by a “relative weight”. In the CLI input the relative weights can be specified using the following set of flags (instead of the “percentChances…” flag):
@@
[…]
@@
Table 1: Input flags to specify the relative weights of various genetic operations, with their default values. The actual percentage chances of applying a genetic operation are listed in the last columns for the VC and FC/MC search types.
@@
It should be noted that our benchmarking tests performed on the Ti-O system show that the number of unique chemical compositions produced in a VC search increases by 30% or more with the use of permutomic and permucomp operations, with the default genetic operation weights.
@@ 5.5 Secondary mutation: random supercell expansion
If this flag is specified to be a non-zero value, the expansion is performed by randomly choosing three integers as multiples for the lattice vectors, such that the produced supercell has up to “maxAtoms” total atoms in the cell. Further, a randomly chosen atom in the supercell is displaced randomly, subject to the minimum interatomic distances settings (Fig. 3). This operation can be useful in ground state searches where charge density waves lower the energy of the system.
@@ 6 Reference energies
As of version 14, XtalOpt uses the distance above the convex hull as the target value for global energy minimization, as first introduced by Zunger and co-workers [44]. The calculation of the convex hull (performed using the Qhull library [67]), requires defining the reference energies. By default, the code uses “0.0” as the “elemental reference energies”. To obtain the correct convex hull for MC and VC searches, however, the correct reference energies need to be specified.
@@
The energies introduced in the input must have the same unit as is used by the local optimizer (e.g. meV, eV or Hartree) for the chemical formula listed to produce the correct results.
@@ 7 Pareto optimization and parent selection
In release 13 of XtalOpt [59] a basic generalized fitness function, calculated as the weighted sum of normalized values of objectives, was introduced to perform a MOGO search for materials with desired functionalities. The effectiveness of this formalism has been demonstrated through powder X-ray diffraction-assisted discovery of meta-stable phases of Ti-O [43], and the search for high-pressure electride phases of Ca5Pb3 using electronic charge density information [68]. Further, it was previously shown that filtering the parent pool based upon structural features that include the local crystalline order and symmetry is an effective way to find metastable phases [65].
By default XtalOpt employs this scalar fitness function for global optimization (i.e., the “Basic” optimization type) regardless of the number of objectives. In the new version of XtalOpt Pareto optimization, which is in principle applicable to both SO and MO searches, has been implemented as an alternative (the “Pareto” optimization type). The user can instruct the code to utilize Pareto optimization through the input flag:
@@
The Pareto-based fitness value relies on the calculated rank of structures, rr, which is a value in the [0,N1][0,N-1] range , where NN is the total number of Pareto fronts and r=0r=0 corresponds to the global Pareto front. Using the obtained ranks, solutions of each front are assigned a raw fitness value of
[…]
@@
[…]
@@
where the default value of “1-1” implies not rounding the objective values.
@@ 9 Similarity check
The first version of XtalOpt used an approximate scheme, identifying duplicates by comparing their spacegroups, volumes and enthalpies [26] to within a user-specified tolerance. Because this method was prone to false positives, an exact comparison technique that searches for transformations mapping atoms from one structure to another was developed. The resulting XtalComp library [70] was used to find duplicates since version 9 [61] of XtalOpt. Duplicates were marked as such in the output and removed from the run. Though this exact matching scheme was key for finding certain inorganic compounds [71], it is dependent upon the choice of the user-specified tolerance and does not work well for molecular systems. For example, it would likely fail to flag structures that differ by a slight rotation of a molecule or a functional group as duplicates because the required tolerance would be too large.
@@
By default, the RDF tolerance is set to zero; hence XtalComp is employed for duplicate checking to maintain consistency with previous versions of XtalOpt. In general, however, the use of RDF check is recommended, as it provides a quantitative measure for the similarity of structures that can be visualized using common plotting software.
@@ 10 Seed structures
One example where providing sub-system seeds could be useful is for obtaining energies of reference structures, to be subsequently used in the convex hull construction, that are consistent with the local optimization settings of the search. This option, however, should be utilized carefully since a failure of the local relaxation of the reference structure by the external code will result in an incorrect convex hull leading to errors in fitness determinations. The introduction of sub-systems may also be useful if they contain certain structural motifs that could be present in the full system.
The seed structures, in general, are not required to comply with the minimum and maximum number of atom settings. Once a seed structure is read in with a total atom count that is not consistent with the existing settings, the atom counts limits will be updated, while a warning message is produced in the output.
@@ 11 Construction of molecular units
For an MC or VC search, it is possible to use molecular units for the construction of the initial random generation of structures, as first introduced in XtalOpt version 10 [75] for a FC search. However, now the number of atoms of each type that can participate in the construction of molecular units is limited to the smallest number of atoms of that type in the input list of chemical formulas, regardless of search type.
@@ 12 Interface to machine-learning interatomic potentials
An easy-to-use Python script, vasp_uip.py, is included in the new release of XtalOpt to facilitate such calculations. This wrapper enables local optimizations with a series of the recently developed universal interatomic potentials (UIPs), MACE [52], CHGNet [53], Orb [54], and MatterSim [55], to perform local structure optimization using the Atomic Simulation Environment (ASE) library [76]. The input/output format is that of the VASP code, i.e., the script reads in a POSCAR file and produces a CONTCAR and a minimal OUTCAR file, which follows the VASP format for the outputted entries.
@@
With the above-mentioned input and output file formats, this script is designed to be used in the XtalOpt runs as a VASP optimizer, allowing for the efficient ground state search for a desired chemical system. It should be noted that the implemented potentials have been chosen as prototypes of the increasingly popular UIPs, and extending the script to support other similar platforms is straightforward.
@@ 13.1 Input entry conventions
In general, and for consistency, the input patterns for the CLI flags are designed with the following conventions:
@@ 13.4 Convex hull snapshots
13.5 Computational scaling
In order to measure the computational cost of various XtalOpt tasks in a typical search, we performed a series of benchmarking test runs on the Ti-O system to generate 100, 400, and 1000 structures. These searches were conducted over a variety of input parameters, e.g., VC and non-VC searches, basic and Pareto optimization, with and without additional objectives, and using RDF and XtalComp similarity checks. The runs were performed on 24 cores with 20 parallel local optimizations, and the vasp_uip.py script with MACE UIP was employed for structural relaxations.
The profiling results were then used to measure the CPU time usage of various XtalOpt functions and tasks, and compare them against the overall XtalOpt CPU usage and the total CPU time of the run, i.e., the total time consumed by the code, external objective calculations, and the local optimizations that correlates with the run wall time.
Based on the above analysis, on average, XtalOpt consumes 35%, 17%, and 10% of the total CPU time in generating 100, 400, and 1000 structures, respectively. This is an expected result, as the increasing number of generated structures in a search requires more CPU time for local optimizations (and objective calculations, if applicable) performed by the external codes. A breakdown of the CPU time usage for various tasks, averaged over all searches of different settings but the same target structure count, is listed in Table 3. The results show that the overall process of the parent selection is the most demanding task in a search, consuming on average $\sim$7of the code’s CPU usage (0.9% of the total CPU time).
Task Share of XtalOpt CPU time Share of total CPU time
Input/Output 1.86% 0.38%
Similarity check 2.53% 0.53%
Fitness analysis and parent selection 6.95% 0.90%
Structure generation and analysis 3.98% 0.77%
Table 3: Relative CPU time usage of various XtalOpt tasks, obtained by averaging the results of multiple profiling runs with different input settings; compared to the code and total CPU time usages.
Further, the most computationally-demanding XtalOpt functions were identified for various search settings. The average share from the code and total CPU times for each function was calculated over all searches with the same total number of generated structures. According to the average CPU time usages, the non-dominated sorting and comparison of RDF vectors were found to be the most computationally expensive functions in a typical search. Currently, and by default, a non-dominated sorting is conducted regardless of the global optimization type to obtain and output the Pareto front indices for user’s information. This operation, on average, consumes 6% of the overall XtalOpt CPU usage (0.8% of the total CPU time) in a typical search, ranking the function as the most computationally expensive task. Second in the CPU usage is, when employed, the comparison of RDF vectors with using 3.7% of the overall XtalOpt CPU time (0.7% of the total CPU time).
It should be noted that the share of computational cost of the code and its functions from the total CPU usage of the runs is measured for searches conducted using fast UIP local optimizations. Therefore, the computational overhead of the XtalOpt tasks would be insignificant when the structural relaxations are performed using ab initio codes.